OpenStudy (anonymous):

solve the following equations: log3 (x+1) = log9 (1-X)

OpenStudy (anonymous):

use change of base formulas

OpenStudy (amistre64):

cross multiply and hope for the best :)

OpenStudy (anonymous):

log {9} (1-x) = (1/2) log {3} (1-x)

OpenStudy (anonymous):

where the number in {} is the base

OpenStudy (amistre64):

yeah, thats prolly base 3 and base 9 there eh

OpenStudy (anonymous):

then you can use normal log laws to collect the logs into a single log

OpenStudy (anonymous):

let y= \[\log_{9} (1-x) \]

OpenStudy (anonymous):

let y= \[\log_{9} (1-x) \]

OpenStudy (anonymous):

so \[9^{y} = 1-x \]

OpenStudy (anonymous):

\[3^{2y} = 1-x \]

OpenStudy (anonymous):

take log base 3 of both sides , so it matches with the other log in the question

OpenStudy (anonymous):

then divide by two , thats how you change the bases

OpenStudy (anonymous):

rest is easy

OpenStudy (anonymous):

think the only solution is x = 0 this equation says \[(x+1)^2=1-x\] lets try this: \[log_3(x+1)=log_9(1-x)=\frac{log_3(1-x)}{log_3(9)}=\frac{log_3(1-x)}{2}\] \[2log_3(x+1)=log_3(1-x)\] \[log_3(x+1)^2=log_3(1-x)\] \[(x+1)^2=1-x\] \[x^2+2x+1=1-x\] \[x^2+3x=0\] \[x(x+3)=0\] \[x=0\]or \[x=-3\] and of course -3 is out because you cannot take the log of a negative number

OpenStudy (anonymous):

thank you soo much! i like the way you explaine things